# Approximation of multipartite quantum states: revised version with new applications

### v1

@@ 5.3 On the energy-constrained versions of $E_{R,\pi}$
  • the conditional entanglement of mutual information in $A_{1}...A_{n}$ ($N^{n}_{n,1}(1,(n+1)/2)$, [31, Section IV-C]).
Note also that for any nonnegative characteristic $f$ of a multipartite quantum system $A_{1}...A_{n}$ not increasing under local channels acting on the subsystems $A_{1}$,…,$A_{m}$ the implications
$f\in L^{m}_{n}(C,D)\;\Rightarrow\;f\circ\Lambda\in L^{m}_{n}(C,D)\quad\textrm{and}\quad f\in\widehat{L}^{m}_{n}(C,D)\;\Rightarrow\;f\circ\Lambda\in\widehat{L}^{m}_{n}(C,D)$
hold for any channel
$\Lambda=\Phi_{1}\otimes\Phi_{2}\otimes\cdots\otimes\Phi_{n},$ (63)
where $\Phi_{k}$ is a channel from the system $A_{k}$ to any system $B_{k}$, $k=\overline{1,n}$ [31, Section V]. [^23] We assume that the function $f$ is defined on the set of states of any $n$-partite system, in particular, the system $B_{1}...B_{n}$.
This implies, for example, that the family of functions $f_{\Lambda}(\rho)=I(B_{1}\!:\!B_{2}|B_{3})_{\Lambda(\rho)}$, where $\Lambda$ runs over the set of all channels of the form (63) with $n=3$, is contained in the class $L^{1}_{3}(2,2)$ (because the function $f(\rho)=I(A_{1}\!:\!A_{2}|A_{3})_{\rho}$ is nonnegative, belongs to the class $L^{1}_{3}(2,2)$ and does not increase under channels acting on $A_{1}$). Hence, this family has a common faithful $II^{1}_{3}$-type continuity bound (in terms of Section 2.3).
I am grateful to A.S.Holevo, G.G.Amosov and E.R.Loubenets for useful discussion. I am also grateful to L.Lami for the collaboration that motivated this research.

### v2

@@ 4.3.3 Other results about the function $E_{I}$
for any $\,m=2,3,...,n$, where $E_{I}$ denotes the conditional entanglement of mutual information of a state of the system $A_{1}...A_{m}$.
Proof. Since $I(A_{1}\!:...:\!A_{n})_{\rho}<+\infty$ by the upper bound (9), the claim of the proposition follows from Lemma 4.1.4 in Section 4.1.4 and Proposition 4.3.2B in Section 4.3.2. $\Box$
5 Concluding remarks
In this article, we describe a method for studying characteristics of multipartite infinite-dimensional quantum systems based on a finite-dimensional approximation of these characteristics. This method can be applied to any characteristic of a multipartite quantum system provided that a uniform continuity bound with the energy-type constraint of one of two types (described in Section 2.3) is constructed for this characteristic (or, at least, the existence of such a continuity bound is proved).
We consider applications of the new approximating technique for analysis of several important characteristics of composite quantum systems for which the methods proposed before can not be applied. In fact, one can use this technique for studying any characteristics from the classes $I^{m}_{n}$ and $II^{m}_{n}$ introduced in Section 2.3 (an incomplete list of such characteristics is presented in the Appendix).
One of the advantages of the proposed technique is its applicability to analysis of the regularized versions of correlation and entanglement measures in composite quantum systems (which are difficult to analyze using standard methods due to the complexity of their definition). So, the results on the properties of the regularized relative entropy of $\pi$-entanglement and the regularized Rains bound obtained in the article are the most nontrivial applications of this technique.
The nonstandard use of the uniform continuity bounds with the energy-type constraints considered in the article shows the necessity to develop new methods for deriving such continuity bounds. They can be used for solving different questions concerning a given characteristic of infinite-dimensional quantum systems not directly related to its uniform continuity.
Appendix: Basic characteristics from the classes $I^{m}_{n}$ and $II^{m}_{n}$
In the presentation of the main results of this article, the classes $I^{m}_{n}$ and $II^{m}_{n}$ introduced in Section 2.3 were significantly used. Below is an (incomplete) list of characteristics of a quantum system $A_{1}...A_{n}$, indicating the concrete classes to which they belong with the justification. [^25] There is a reason to believe that many of the characteristics marked as belonging to the class $II^{m}_{n}$ actually belong to the class $I^{m}_{n}$. But to proof this conjecture a rigorous proof is needed. For example, it is pointed below that the $n$-partite squashed entanglement and c-squashed entanglement belong to the class $II^{n-1}_{n}$, while the bipartite squashed entanglement is marked as a characteristic from $I^{1}_{2}$, because a $I^{1}_{2}$-type faithful uniform continuity bound for the latter is constructed in [24]. It is natural to assume that updating the arguments in [24] it is possible to prove the existence of $I^{n-1}_{n}$-type faithful uniform continuity bounds for the $n$-partite squashed entanglement and c-squashed entanglement.
Class $I_{n}^{m}$
  • the von Neumann entropy $S(\rho)$ of a state $\rho$ of $A_{1}$ ($I_{1}^{1}$, Section 2.3);
  • the (extended) quantum conditional entropy $S(A_{1}|A_{2})$ defined in (7) ($I_{2}^{1}$, Section 2.3);
  • the quantum mutual information $I(A_{1}\!:...:\!A_{n})$ ($I^{n-1}_{n}$, Section 2.3);
  • the (extended) quantum conditional mutual information $I(A_{1}\!:...:\!A_{n}|A_{n+1})$ ($I^{n-1}_{n+1}$, Section 2.3);
  • the one-way classical correlation $C_{A_{2}}$ in $A_{1}A_{2}$ with measured system $A_{2}$ ($I^{1}_{2}$, Section 2.3));
  • the regularized one-way classical correlation $C^{\infty}_{A_{2}}$ in $A_{1}A_{2}$ with measured system $A_{2}$ ($I^{1}_{2}$, Section 2.3);
  • the quantum discord in $A_{1}A_{2}$ with measured system $A_{2}$ ($I^{1}_{2}$, Section 2.3);
  • the entanglement of formation $E_{F}$ in $A_{1}A_{2}$ ($I^{1}_{2}$, Section 2.3);
  • the regularized entanglement of formation $E^{\infty}_{F}$ in $A_{1}A_{2}$ ($I^{1}_{2}$, Section 2.3);
  • the squashed entanglement $E_{\rm sq}$ in $A_{1}A_{2}$ ($I^{1}_{2}$, Section 2.3);
  • the topological entanglement entropy in $A_{1}A_{2}A_{3}$ ($I_{3}^{1}$, the representation in [16, Remark 9] and the properties of the QCMI mentioned in [26, Section 4.2]).
Class $II_{n}^{m}$
  • all the characteristics belonging to the class $I_{n}^{m}$ (described before);
  • the one-way classical correlation $C_{A_{1}}$ in $A_{1}A_{2}$ with measured system $A_{1}$ ($II^{1}_{2}$, Section 2.3);
  • the regularized one-way classical correlation $C^{\infty}_{A_{1}}$ in $A_{1}A_{2}$ with measured system $A_{1}$ ($II^{1}_{2}$, Section 2.3);
  • the quantum discord in $A_{1}A_{2}$ with measured system $A_{1}$ ($II^{1}_{2}$, Section 2.3);
  • the constrained Holevo capacity $\chi_{A_{1}}$ of the partial trace over $A_{2}$ in $A_{1}A_{2}$ ($II^{1}_{2}$, [26, Proposition 21]);
  • the relative entropy of $\pi$-entanglement $E_{R,\pi}$ in $A_{1}...A_{n}$ for any set of partitions $\pi$ ($II^{n-1}_{n}$, Section 2.3);
  • the regularized relative entropy of $\pi$-entanglement $E^{\infty}_{R,\pi}$ in $A_{1}...A_{n}$ for any singleton set of partitions $\pi$ ($II^{n-1}_{n}$, Section 2.3);
  • the relative entropy of NPT-entanglement $E_{R,PPT}^{\infty}$ in $A_{1}A_{2}$ ($II^{1}_{2}$, Remark 4.1.3);
  • the regularized relative entropy of NPT-entanglement $E_{R,PPT}^{\infty}$ in $A_{1}A_{2}$ ($II^{1}_{2}$, Remark 4.1.3);
  • the Rains bound in $A_{1}A_{2}$ ($II^{1}_{2}$, Section 4.22);
  • the regularized Rains bound in $A_{1}A_{2}$ ($II^{1}_{2}$, Section 4.22);
  • the squashed entanglement $E_{\rm sq}$ in $A_{1}...A_{n}$ ($II^{n-1}_{n}$, [26, Proposition 24]);
  • the c-squashed entanglement $E_{\rm c\textup{-}sq}$ in $A_{1}...A_{n}$ ($II^{n-1}_{n}$, [26, Proposition 24]);
  • the entropy reduction of a local measurement in $A_{1}A_{2}$ with measured system $A_{1}$ ($II^{1}_{2}$, [63, Corollary 3]);
  • the function $\rho\mapsto I(B_{1}\!:...:\!B_{n})_{\Phi_{1}\otimes...\otimes\Phi_{n}(\rho)}$ on $\mathfrak{S}(\mathcal{H}_{A_{1}...A_{n}})$, where $\Phi_{1}:A_{1}\rightarrow B_{1}$,…, $\Phi_{n}:A_{n}\rightarrow B_{n}$ are arbitrary quantum channels ($II^{n-1}_{n}$, [15, Proposition 24]). [^26] The observation in [15, Section V] shows that many other functions of this type belong to the class $II^{m}_{n}$.
I am grateful to A.S.Holevo, G.G.Amosov and E.R.Loubenets for useful discussion. I am grateful to M.Plenio and to the participants of his seminar at Ulm University (where the initial version of this work was reported) for useful comments. I am also grateful to L.Lami for the collaboration that motivated the final stage of this research. Special thanks to S.Becker for construction of a counterexample showing that the condition (19) is not a criterion of the FA-property.